OpenStudy (anonymous):

The difference of 2consecutive numbers is an odd number.true of false?

OpenStudy (kirbykirby):

3-2=? 5-4 = ? 7-8 = ? 100-99 = ?

OpenStudy (kirbykirby):

More generally, if x is a number, then a consecutive number is x+1 \(x - (x+1)=?\) you can also do : \((x+1)-x = ?\)